QuestionCoveBot:

When x=2 and y=-2 What is t? x(7-12y)*16-x4(6+y) = t

QuestionCoveBot:

\[x(7-12y)*16-x4(6+y) = t\]

KynxKatz:

Let's solve for t. x(7−12y)(16)−x4(6+y)=t Step 1: Flip the equation. t=−x4y−6x4−192xy+112x Answer: t=−x4y−6x4−192xy+112x

QuestionCoveBot:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @KynxKatz Let's solve for t. x(7−12y)(16)−x4(6+y)=t Step 1: Flip the equation. t=−x4y−6x4−192xy+112x Answer: t=−x4y−6x4−192xy+112x \(\color{#0cbb34}{\text{End of Quote}}\) Simplified form please?

KynxKatz:

Sure

KynxKatz:

Simplified answer is: t=112x−192xy−6x4−x4y

KynxKatz:

Hats the closest answer i can give u

KynxKatz:

Thats*

QuestionCoveBot:

I'm still confused about this... @darkknight

boredfr:

Ok so if you replace x with 2 and y with negative 2 then your equation should look like this: 2[(7-12)(-2)]*16-(2)4(6+-2)=t Do you know what the next step is?

QuestionCoveBot:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @boredfr Ok so if you replace x with 2 and y with negative 2 then your equation should look like this: 2[(7-12)(-2)]*16-(2)4(6+-2)=t Do you know what the next step is? \(\color{#0cbb34}{\text{End of Quote}}\) No.

boredfr:

You are going to simplify both sides of the equation Do you know how to set that up?

QuestionCoveBot:

Uh no. I am just straight up confused with this. I was once smart.

boredfr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @QuestionCoveBot Uh no. I am just straight up confused with this. I was once smart. \(\color{#0cbb34}{\text{End of Quote}}\) Lmao it's ok I got you, just gimme a minute to write it out

boredfr:



KynxKatz:

How did u get 2 n those things

boredfr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @KynxKatz How did u get 2 n those things \(\color{#0cbb34}{\text{End of Quote}}\) What 2 things?

QuestionCoveBot:

So the answer is t= 256?

KynxKatz:

Nvm

boredfr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @QuestionCoveBot So the answer is t= 256? \(\color{#0cbb34}{\text{End of Quote}}\) Yes but do you understand how to get the answer?

QuestionCoveBot:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @boredfr \(\color{#0cbb34}{\text{Originally Posted by}}\) @QuestionCoveBot So the answer is t= 256? \(\color{#0cbb34}{\text{End of Quote}}\) Yes but do you understand how to get the answer? \(\color{#0cbb34}{\text{End of Quote}}\) A little.

boredfr:

What is it that you don't understand?

QuestionCoveBot:

Ooh, wait a minute. Now I understand. Took me a bit. ._. Thanks.

boredfr:

Np :P